# Parallel Split Learning with Global Sampling

### v4

@@ Abstract
Distributed deep learning in resource-constrained environments faces scalability and generalization challenges due to large effective batch sizes and non-identically distributed client data. We introduce a server-driven sampling strategy that maintains a fixed global batch size by dynamically adjusting client-side batch sizes. This decouples the effective batch size from the number of participating devices and ensures that global batches better reflect the overall data distribution. Using standard concentration bounds, we establish tighter deviation guarantees compared to existing approaches. Empirical results on a benchmark dataset confirm that the proposed method improves model accuracy, training efficiency, and convergence stability, offering a scalable solution for learning at the network edge.
Index Terms:
Parallel split learning, distributed deep learning, non-iid data, IoT.
@@ I Introduction
To address training latency in SL, parallel split learning (PSL) [8, 9, 10, 11, 12] has been proposed and introduces parallelism without the need for multiple server model instances like in split federated learning (v1) [13] or client model aggregation like in FedAvg [14]. However, PSL encounters the large effective batch size problem [15, 11], where the effective batch size scales with the number of clients. Additionally, devices usually contribute from local datasets that are not independent and identically distributed (IID) and vary in sizes, causing batch distributions to deviate from the overall data distribution, which can lead to poor generalization. This issue of non-IID data is a general problem in DDL, not specific to PSL.
Figure 1: Schematic illustration of the proposed GPSL method with variable local batch sizes summing to a fixed global batch size, mitigating the large effective batch size and non-IID data problems in PSL.
We propose parallel split learning with global sampling (GPSL) building upon the original PSL framework introduced in [8] to address the large effective batch size problem and the issue of non-IID data. GPSL allows for variable local batch sizes that sum up to a fixed global batch size. This decouples the number of clients from the effective batch size. We show that this decoupling is crucial for controlling the deviation probability of global batches from the overall data distribution, leading to improved model performance and reduced training latency in non-IID settings. In particular, we show that the choice of local batch sizes is crucial for controlling the deviation probability. An overview of GPSL is shown in Figure 1. The main contributions of this paper are listed below:
  • We introduce the GPSL method to address the large effective batch size problem and the non-IID data problem in PSL.
  • We analyze the global batch deviation and establish tight deviation bounds for GPSL.
  • Our simulation results show that GPSL reduces the deviation probability of global batches significantly, leading to improved accuracy and reducing training latency.
@@ II Related Work
Jeon and Kim [8] introduce PSL and use gradient averaging over all clients to address the client decoupling problem, where the information flow between clients is decoupled during BP. They fixed the local batch sizes proportional to the dataset size of each client. Lyu et al. [17] propose averaging the clients’ local loss function outputs instead of their local gradients. This approach aligns with that of [8] due to the linearity of the gradient operator, but enhances privacy as clients can compute the local loss functions without sharing labels. Both show stable convergence and improved accuracy in both IID and non-IID settings. However, the proportion of each client’s data in the total was chosen to yield an integer when scaled by the global batch size to determine local batch sizes. In practice, rounding is typically necessary when determining each client’s local batch size, which can lead to imbalances in the global batch composition. Furthermore, the large effective batch size problem is not addressed in these works.
@@
Cai and Wei [20] propose an incentive mechanism to handle non-IID data in PSL, which encourages clients with larger and more evenly distributed datasets to participate in training. The bargaining game approach for personalized SL [21] balances energy consumption and privacy in non-IID settings. However, these methods may face scalability issues and may not fully address the heterogeneity of data distributions across diverse clients.
@@ III Preliminaries
We consider the PSL setting from [8]. The client index set is $\mathcal{K}=\{1,2,\dots,K\}$. A server and $|\mathcal{K}|=K$ clients train a neural network model with parameter vector $\boldsymbol{w}$ collaboratively using the PSL framework. We focus on supervised classification tasks, where the dataset $\mathcal{D}_{k}$ of each client $k\in\mathcal{K}$ consists of input-label pairs $\mathcal{D}_{k}=\{(\boldsymbol{x}_{k,i},y_{k,i})\}_{i=1}^{D_{k}}$, where $D_{k}\in\mathbb{N}$ is the dataset size, $\boldsymbol{x}_{k,i}$ is the input data, and $y_{k,i}\in\mathcal{M}$ is the ground truth class label of $\boldsymbol{x}_{k,i}$. Here, =$\mathcal{M}=\{1,2,\dots,M\}$ represents all classes and $M\in\mathbb{N}$ is the number of classes. The overall dataset $\mathcal{D}_{0}=\bigcup_{k\in\mathcal{K}}\mathcal{D}_{k}$ is the union of all clients’ datasets, with a total size $D_{0}=\sum_{k\in\mathcal{K}}D_{k}$.
@@
where $c(\boldsymbol{y},m)$ counts the number of times class $m$ appears in $\boldsymbol{y}$. The overall class distribution of $\mathcal{D}_{0}$ with class label vector $\boldsymbol{y}_{0}\in\mathcal{M}^{D_{0}}$ is denoted as $\boldsymbol{\beta}_{0}\in\mathbb{R}^{M}$ according to the same definition.
Every epoch consists of $T\in\mathbb{N}$ optimization steps. The set of optimization steps is denoted as $\mathcal{T}=\{1,2,\ldots,T\}$. Each client $k$ has a local batch size $B_{k}^{(t)}\in\mathbb{N}$. The global batch size $B^{(t)}=\sum_{k\in\mathcal{K}}B_{k}^{(t)}$ is the total number of intermediate representations that arrive at the server in each optimization step $t\in\mathcal{T}$. Accordingly, we differentiate between local batches $\mathcal{B}^{(t)}_{k}\subseteq\mathcal{D}_{k}$ and the global batch $\mathcal{B}^{(t)}=\bigcup_{k\in\mathcal{K}}\mathcal{B}^{(t)}_{k}$.
@@
    Server-Side FP: The server concatenates the received activations to obtain the global batch ${\mathcal{B}^{(t)}=\bigcup_{k\in\mathcal{K}}f(\mathcal{B}^{(t)}_{k};\boldsymbol{w}_{c,k})}$ and computes the final predictions $f_{s}(\mathcal{B}^{(t)};\boldsymbol{w}_{s})$ to compute the average loss.
@@ IV System Model
This section introduces the proposed GPSL method and analyzes how it improves the alignment between global batches and the overall data distribution compared to random uniform sampling in a centralized setting.
@@ IV-A Parallel Split Learning with Global Sampling
In GPSL, we replace the notion of fixed per‐client batch sizes with a single global batch size $B$. The server first gathers each client’s dataset size $D_{k}$. This information can trivially be obtained from the server through counting the number of samples that arrive from each client and is not sensitive. Before training the server must choose how many samples $B_{k}^{(t)}$ each client $k$ will contribute to each training step $t\in\mathcal{T}$, subject to
@@
Rather than uniformly fixing these $B_{k}^{(t)}$, we draw the local batch sizes using a global sampling procedure over the pooled data. We first set each client’s remaining dataset size $D_{k}^{(0)}=D_{k}$. Then, for each $t\in\mathcal{T}$:
@@
      $\pi_{k}^{(i)}=D_{k}^{(i)}/\sum_{j\in\mathcal{K}}D_{j}^{(i)}.$ (2)
    2. (b)
      Sample a client index $z_{i}\sim\mathrm{Categorical}(\boldsymbol{\pi}^{(i)})$.
    3. (c)
      Decrement its remaining dataset size
      $D_{z_{i}}^{(i+1)}\leftarrow D_{z_{i}}^{(i)}-1.$ (3)
@@
    Batch Composition. After $B$ draws, the number of samples from client $k$ are obtained by
    $B_{k}^{(t)}\;=\;|\{\,i:z_{i}=k\}|.$ (4)
    Remaining dataset sizes are reset to $D_{k}^{(0)}=D_{k}^{(i)}$
The server forwards the $\{B_{k}^{(t)}\}^{t\in\mathcal{T}}$ to each client before training starts. During training clients uniformly sample random mini-batches using $B_{k}^{(t)}$ as local batch size. Because each draw in step 1) selects one of the remaining points uniformly without replacement, the resulting batch during training is exactly a uniform draw at random of $B$ distinct samples from the union $\mathcal{D}_{0}$. The rest of the PSL training loop remains unchanged, allowing seamless integration of the sampling procedure into existing PSL frameworks.
@@ IV-B Deviation Analysis
We assess how well a batch $\mathcal{B}\subset\mathcal{D}_{0}$ of size $B$ reflects the true label distribution $\boldsymbol{\beta}_{0}$ using the $\ell_{1}$ deviation:
@@
[…]
@@
so that all finite‐population corrections remain positive and $\varepsilon B\leq B-1$. When $\mathcal{B}$ is drawn by sampling $B$ points without replacement from $\mathcal{D}_{0}$, the count $Y_{m}=c(\mathcal{B},m)$ follows a hypergeometric distribution
@@
$\Pr\bigl{(}|Y_{m}-B\beta_{0,m}|\geq\varepsilon B\bigr{)}$ $=\Pr\!\Bigl{(}\bigl{|}\tfrac{Y_{m}}{B}-\beta_{0,m}\bigr{|}\geq\varepsilon\Bigr{)}$ (7)

### v5

@@ Abstract
Parallel split learning (PSL) suffers from two intertwined issues: the effective batch size grows with the number of clients, and data that is not identically and independently distributed (non-IID) skews global batches. We present parallel split learning with global sampling (GPSL), a server-driven scheme that fixes the global batch size while computing per-client batch-size schedules using pooled-level proportions. The actual samples are drawn locally without replacement by each selected client. This eliminates per-class rounding, decouples the effective batch from the client count, and makes each global batch distributionally equivalent to centralized uniform sampling without replacement. Consequently, we obtain finite-population deviation guarantees via Serfling’s inequality, yielding a zero rounding bias compared to local sampling schemes. GPSL is a drop-in replacement for PSL with negligible overhead and scales to large client populations. In extensive experiments on CIFAR-10/100 and ResNet-18/34 under non-IID splits, GPSL stabilizes optimization and achieves centralized-like accuracy, while fixed local batching trails by up to 60%. Furthermore, GPSL shortens training time by avoiding inflation of training steps induced by data-depletion. These findings suggest GPSL is a promising and scalable approach for learning in resource-constrained environments.
@@ I Introduction
To address training latency in SL, parallel split learning (PSL) [8, 9, 10, 11, 12] has been proposed and introduces parallelism without the need for multiple server model instances like in split federated learning (v1) [13] or client model aggregation like in FedAvg [14]. However, PSL encounters the large effective batch size problem [15, 11]. As the client count grows, per-client batching with fixed local batch sizes makes the effective global batch size scale with the number of clients. This reduces gradient noise but can hurt generalization unless hyperparameters are carefully retuned [16, 17]. It also increases server memory footprint and per-step latency.
Additionally, devices usually contribute from local datasets that are not independent and identically distributed (IID) and vary in sizes. This issue of non-IID data is a general problem in DDL, not specific to PSL. Under non-IID data, local sampling schemes skew global batch composition due to rounding and per-client constraints, inflating deviation from the pooled distribution and slowing or destabilizing convergence. Rounding also accelerates client data depletion, increasing the number of steps per epoch and prolonging training.
Figure 1: GPSL schematic: variable local batch sizes sum to a fixed global batch size, mitigating large effective batch size and non-IID issues in PSL.
We propose parallel split learning with global sampling (GPSL) building upon the original PSL framework introduced in [8] to address these problems. It fixes a global batch size and assigns per-client batch-size schedules derived from pooled-level proportions using only dataset-size metadata. Each client then samples its own data locally and without replacement to meet its scheduled count. This suppresses per-class rounding and decouples the effective batch size from the number of clients. Under GPSL, each global batch closely matches the pooled distribution in expectation and inherits finite-population deviation guarantees with zero rounding bias compared to local sampling schemes.
GPSL remains a drop-in replacement for PSL with negligible overhead. This means it integrates with existing PSL components such as client selection, clustering, and resource-aware schedulers [18, 10, 15]. With a fixed global batch size and global sampling, it scales to large client populations without increasing server memory or per-step latency, and remains stable under non-IID splits. This makes it practical for edge learning: it requires only dataset-size metadata, uses lightweight client models with a single shared server model, and adds minimal coordination overhead.
An overview of GPSL is shown in Figure 1. The main contributions of this paper are:
  • Novel sampling mechanism: A server-driven global sampling scheme that fixes the global batch size while allocating dynamic local batch sizes, decoupling the effective batch size from the number of clients and eliminating per-class rounding effects present in local sampling schemes.
  • Finite-population guarantees: We derive deviation bounds via Serfling’s inequality with the finite-population correction, yielding zero rounding bias compared to local sampling schemes and showing that GPSL is equivalent to centralized uniform sampling without replacement.
@@ II Related Work
Jeon and Kim [8] introduce PSL and use gradient averaging over all clients to address the client decoupling problem, where the information flow between clients is decoupled during back propagation (BP). They fixed the local batch sizes proportional to the dataset size of each client. Lyu et al. [20] propose averaging the clients’ local loss function outputs instead of their local gradients. This approach aligns with that of [8] due to the linearity of the gradient operator, but enhances privacy as clients can compute the local loss functions without sharing labels. Both show stable convergence and improved accuracy in both IID and non-IID settings. However, the proportion of each client’s data in the total was chosen to yield an integer when scaled by the global batch size to determine local batch sizes. In practice, rounding is typically necessary when determining each client’s local batch size, which can lead to imbalances in the global batch composition. Furthermore, the large effective batch size problem is not addressed in these works.
@@
Cai and Wei [21] propose an incentive mechanism to handle non-IID data in PSL, which encourages clients with larger and more evenly distributed datasets to participate in training. However, it introduces coordination and participation-scoring overhead and relies on stable client availability. Other efforts to improve PSL focus on reducing training and communication latency [9, 10, 22]. Methods include last-layer gradient aggregation, resource optimization, client clustering, and cut-layer selection.
@@ III Preliminaries
Before introducing our GPSL method, we first present the system model and preliminaries of the PSL framework from [8] that we build upon. The client index set is $\mathcal{K}=\{1,2,\dots,K\}$. A server and $|\mathcal{K}|=K$ clients train a neural network model with parameter vector $\boldsymbol{w}$ collaboratively using the PSL framework. We focus on supervised classification tasks, where the dataset $\mathcal{D}_{k}$ of each client $k\in\mathcal{K}$ consists of input-label pairs $\mathcal{D}_{k}=\{(\boldsymbol{x}_{k,i},y_{k,i})\}_{i=1}^{D_{k}}$, where $D_{k}\in\mathbb{N}$ is the dataset size, $\boldsymbol{x}_{k,i}$ is the input data, and $y_{k,i}\in\mathcal{M}$ is the ground truth class label of $\boldsymbol{x}_{k,i}$. Here, $\mathcal{M}=\{1,2,\dots,M\}$ represents all classes and $M\in\mathbb{N}$ is the number of classes. The pooled dataset $\mathcal{D}_{0}=\bigcup_{k\in\mathcal{K}}\mathcal{D}_{k}$ is the union of all clients’ datasets, with a total size $D_{0}=\sum_{k\in\mathcal{K}}D_{k}$.
@@
where $c(\boldsymbol{y},m)$ counts the number of times class $m$ appears in $\boldsymbol{y}$. The pooled class distribution of $\mathcal{D}_{0}$ with class label vector $\boldsymbol{y}_{0}\in\mathcal{M}^{D_{0}}$ is denoted as $\boldsymbol{\beta}_{0}\in\mathbb{R}^{M}$ according to the same definition.
Every epoch consists of $T\in\mathbb{N}$ optimization steps. The set of optimization steps is denoted as $\mathcal{T}=\{1,2,\ldots,T\}$. Each client $k$ has a local batch size $B_{k}^{(t)}\in\mathbb{N}$. The global batch size $B^{(t)}=\sum_{k\in\mathcal{K}}B_{k}^{(t)}$ is the total number of intermediate representations that arrive at the server in each optimization step $t\in\mathcal{T}$. Accordingly, we differentiate between local batches $\mathcal{B}^{(t)}_{k}\subseteq\mathcal{D}_{k}$ and global batches $\mathcal{B}^{(t)}=\bigcup_{k\in\mathcal{K}}\mathcal{B}^{(t)}_{k}$.
@@
    Server-Side FP: The server concatenates the received intermediate representations to obtain the global batch $\mathcal{B}^{(t)}=\bigcup_{k\in\mathcal{K}}f(\mathcal{B}^{(t)}_{k};\boldsymbol{w}_{c,k})$} and computes the final predictions $f_{s}(\mathcal{B}^{(t)};\boldsymbol{w}_{s})$ to compute the average loss.
@@
We adopt a generalized PSL formulation that permits variable local batch sizes across $k$ and $t$ to ensure compatibility with the proposed method. In standard PSL, by contrast, the local batch size is typically fixed over time, i.e., $B_{k}^{(t)}=B_{k}$ for all $t\in\mathcal{T}$ [8, 20, 15]. Likewise, if the global batch size is fixed, we write $B^{(t)}=B$. When the per-client sequence $\{B_{k}^{(t)}\}_{t\in\mathcal{T}}$ varies with $t$, we refer to it as the per‑client batch‑size schedule or simply schedule.
@@ IV System Model
This section introduces the proposed GPSL method and analyzes how it improves the alignment between global batches and the pooled data distribution compared to random uniform sampling in a centralized setting.
@@ IV-A Parallel Split Learning with Global Sampling
In GPSL, we replace the notion of fixed per‐client batch sizes $B_{k}$ with per-client schedules $\{B_{k}^{(t)}\}_{t\in\mathcal{T}}$ that sum across clients to a fixed global batch size $B$ in each step $t$. The server first gathers each client’s dataset size $D_{k}$. This information could trivially be obtained from the server through counting the number of samples that arrive from each client and is not sensitive. Before training the server must choose how many samples $B_{k}^{(t)}$ each client $k$ will contribute to each training step $t\in\mathcal{T}$, subject to
@@
Rather than uniformly fixing these $B_{k}^{(t)}$, the server computes local batch sizes by sampling client indices using the pooled proportions implied by the remaining counts, i.e., the number of unused samples per-client. No raw data is accessed by the server and clients later draw their own examples locally to meet the assigned counts. We initialize the number of remaining unused datapoints per-client as $R_{k}\leftarrow D_{k}$ for all $k\in\mathcal{K}$. For each step $t\in\mathcal{T}$, the server allocates the per-step schedule $\{B_{k}^{(t)}\}_{k\in\mathcal{K}}$ across clients according to the pooled proportions using only the remaining counts $\{R_{k}\}_{k\in\mathcal{K}}$:
For $i=1,\dots,B$:
@@
    $\pi_{k}=\frac{R_{k}}{\sum_{j\in\mathcal{K}}R_{j}}.$ (2)
@@
    Sample a client index
    $z\sim\mathrm{Categorical}(\pi_{1},\dots,\pi_{K}).$
  3. 3.
    Increment the scheduled count $B_{z}^{(t)}\leftarrow B_{z}^{(t)}+1$.
  4. 4.
    Decrement the remaining count $R_{z}\leftarrow R_{z}-1$.
The server precomputes and broadcasts the schedules $\{B_{k}^{(t)}\}_{t\in\mathcal{T}}$ to all clients. At step $t$, client $k$ samples $B_{k}^{(t)}$ previously unused examples uniformly without replacement from $\mathcal{D}_{k}$ and then performs the regular PSL steps (see Section III). A final partial batch (if any) may be kept or skipped. No other part of the PSL loop changes, so this sampling procedure drops into existing PSL implementations.
The full procedure (including handling dataset depletion across optimization steps) is summarized in Algorithm 1. The time complexity is $O(BT)$, identical (up to constants) to fixed local batching.
Algorithm 1 Global Sampling Procedure for GPSL
1:Number of clients $K$; client dataset sizes $D_{1},\dots,D_{K}$; global batch size $B$.
2:Local schedules $\{B_{k}^{(t)}\}_{t\in\mathcal{T}}$ for each $k\in\mathcal{K}$.
[…]
5:while $\sum_{k\in\mathcal{K}}R_{k}>0$ do
6:  $B_{k}^{(t)}\leftarrow 0$ for all $k\in\mathcal{K}$ $\triangleright$ Initialize scheduled counts
7:  $L\leftarrow\min\big(B,\sum_{j=1}^{K}R_{j}\big)$ $\triangleright$ Number of draws this step
8:  for $i=1$ to $L$ do
9:   $\pi_{k}\leftarrow R_{k}/\sum_{j=1}^{K}R_{j}$ for all $k\in\mathcal{K}$ with $R_{k}>0$.
10:   $z\sim\mathrm{Categorical}(\pi_{1},\dots,\pi_{K})$.
11:   $B_{z}^{(t)}\leftarrow B_{z}^{(t)}+1$.
12:   $R_{z}\leftarrow R_{z}-1$.
13:  end for
14:  $t\leftarrow t+1$
15:end while
16:Send $\{B_{k}^{(t)}\}_{k\in\mathcal{K}}$ to all clients.
@@ IV-B Deviation Analysis
We assess how well a batch $\mathcal{B}\subset\mathcal{D}_{0}$ of size $B$ reflects the true pooled class distribution $\boldsymbol{\beta}_{0}$ using the $\ell_{1}$ deviation:
@@
where $c(\mathcal{B},m)$ is the count of class-$m$ samples in $\mathcal{B}$. Here, $\boldsymbol{\beta}_{0}$ denotes the pooled class distribution defined in Section III and $D_{0}=\sum_{k\in\mathcal{K}}D_{k}$ is the total number of samples across all clients. Throughout we assume
@@
so that: (i) the finite–population correction term $1-\tfrac{B-1}{D_{0}}>0$ appearing in Serfling’s inequality remains positive (this requires $B\leq D_{0}$); (ii) the deviation event $\lvert Y_{m}-B\beta_{0,m}\rvert\geq\varepsilon B$ is non-vacuous because the maximal attainable deviation for a count bounded in $[0,B]$ is strictly less than $B$ (hence we enforce $\varepsilon B\leq B-1$); and (iii) choosing $\varepsilon>1-1/D_{0}$ would correspond to ruling on deviations that cannot occur under a hypergeometric draw (since at least one sample of any present class must exist in $\mathcal{D}_{0}$). These mild constraints avoid degenerate parameter settings and keep the bound informative. When $\mathcal{B}$ is drawn by sampling $B$ points without replacement from $\mathcal{D}_{0}$, the count $Y_{m}=c(\mathcal{B},m)$ follows a hypergeometric distribution
@@
$\mathrm{P}\bigl(|Y_{m}-B\beta_{0,m}|\geq\varepsilon B\bigr)$ $=\mathrm{P}\!\Bigl(\bigl|\tfrac{Y_{m}}{B}-\beta_{0,m}\bigr|\geq\varepsilon\Bigr)$ (5)
$\leq 2\exp\!\Bigl(-\frac{2\,\varepsilon^{2}\,B}{\,1-\tfrac{B-1}{D_{0}}\,}\Bigr).$ (6)
Although the $\{Y_{m}\}_{m\in\mathcal{M}}$ are dependent as they are required to sum up to $B$, if $\Delta(\mathcal{B},\boldsymbol{\beta}_{0})=\sum_{m\in\mathcal{M}}|Y_{m}/B-\beta_{0,m}|\geq M\varepsilon$, then for some $m$ we have $|Y_{m}/B-\beta_{0,m}|\geq\varepsilon$. A union bound over $m$ yields
$\mathrm{P}\!\Bigl(\Delta(\mathcal{B},\boldsymbol{\beta}_{0})\geq M\varepsilon\Bigr)$ $\leq\sum_{m\in\mathcal{M}}2\exp\!\Bigl(-\tfrac{2\,\varepsilon^{2}\,B}{1-\tfrac{B-1}{D_{0}}}\Bigr)$ (7)
$=2M\exp\!\Bigl(-\tfrac{2\,\varepsilon^{2}\,B}{1-\tfrac{B-1}{D_{0}}}\Bigr).$ (8)
In Algorithm 1, at each draw the server samples a client index $k$ with probability $\pi_{k}=R_{k}/\sum_{j\in\mathcal{K}}R_{j}$, conceptually decrements one remaining item for that client, and updates the counts $\{R_{k}\}_{k\in\mathcal{K}}$. Conditional on the current $\{R_{k}\}_{k\in\mathcal{K}}$, this index-sampling is analytically equivalent to selecting one element uniformly at random from the pooled set of unused samples, even though the server never accesses raw examples. Clients later draw their own data to fulfill the assigned counts. Iterating this for $B$ draws yields a uniform random sample of size $B$ without replacement at the pooled level. Consequently, for any class $m$, the batch count $Y_{m}=c(\mathcal{B},m)$ under GPSL is hypergeometric with the same parameters as centralized sampling, with $\mathbb{E}[Y_{m}]=B\,\beta_{0,m}$. The joint vector $(Y_{m})_{m\in\mathcal{M}}$ thus depends only on $(D_{0},\boldsymbol{\beta}_{0})$ and not on the client partition $\{D_{k}\}_{k\in\mathcal{K}}$ or its skew. It follows that the finite-population Serfling bound in 8 applies to GPSL verbatim.
IV-C Comparison with Fixed Local Batches
Consider fixed local batch sizes $\tilde{B}_{k}=\lceil B\,D_{k}/D_{0}\rceil$ sampled without replacement per client [8]. Let $\tilde{Y}^{\prime}_{m}=\sum_{k\in\mathcal{K}}\tilde{Y}_{k,m}$ with
and define
$\tilde{p}_{m}\coloneqq\sum_{k\in\mathcal{K}}\frac{\tilde{B}_{k}}{B}\,\beta_{k,m}.$
The triangle inequality gives
The sampling term obeys the same Serfling tail as above (conditioned on $\{\tilde{B}_{k}\}_{k\in\mathcal{K}}$). For the deterministic rounding term, since $0\leq\beta_{k,m}\leq 1$,
$\bigl|\tilde{p}_{m}-\beta_{0,m}\bigr|=\Bigl|\sum_{k\in\mathcal{K}}\Bigl(\tfrac{\tilde{B}_{k}}{B}-\tfrac{D_{k}}{D_{0}}\Bigr)\beta_{k,m}\Bigr|\leq\sum_{k\in\mathcal{K}}\tfrac{1}{B}=\tfrac{K}{B}.$ (10)
Let $\delta\coloneqq\max_{m\in\mathcal{M}}|\tilde{p}_{m}-\beta_{0,m}|\leq\tfrac{K}{B}$. Combining 910 and union bounding over $m$ gives the shifted tail bound
$\mathrm{P}\!\Bigl(\Delta(\mathcal{B},\boldsymbol{\beta}_{0})\geq M\varepsilon\Bigr)\leq 2M\exp\!\Bigl(-\tfrac{2\,(\varepsilon-\delta)^{2}\,B}{\,1-\tfrac{B-1}{D_{0}}\,}\Bigr).$ (11)
Hence concentration controls only the random sampling fluctuations. Rounding introduces a fixed bias $\delta$ that does not decay exponentially and can dominate when $K/B$ is not negligible. GPSL avoids this bias ($\delta=0$) while retaining the same exponential decay as 8. Fixed local batches may still perform adequately when $K\ll B$ and rounding effects are negligible. However, such settings are rare in edge learning environments, where large $K$ and small batch sizes dominate due to memory and communication constraints. Moreover, the rounding exacerbates the issue of client data depletion, where client datasets are depleted disproportionately and faster. This increases the total number of training steps $T$ and prolongs training. By contrast, GPSL prevents unnecessary inflation of $T$ by keeping the global batch size $B$ constant, regardless of the number of clients.
V Simulation Results
In this section, we empirically evaluate the performance of GPSL compared to existing sampling methods in PSL. We examine the impact of sampling on batch deviation, training latency, and classification accuracy under both IID and non-IID data splits.
V-A Experimental Setup
We use the CIFAR-10 dataset [24], which consists of $60,000$ color images in $M=10$ classes. The dataset is split into $D_{0}=50,000$ training and $10,000$ test images.